2. Valle de Estós

’PERDIGUERO PEAK 5 hours 3,221 m. high

This peak is where the three big mountain masses of the Ball de Benàs (Maladetas, Posets and Perdiguero) culminate. It is of easy ascent, but the use of the ice axe and crampons are necessary.

Take the forest trail (closed to traffic) of the Ball d’Estós, which will lead you from the ravines up to the Cabana de Sta. Ana (Sta. Ana Hut) in about half an hour. A bit later a path crosses the Estós river through a large wood bridge and begins ascending the opposite slope to the NW, to the La Coma refuge (1,850 m.). You will get there in about 1 hour and 15 minutes. Keep going up by the open ravine of the Perdiguero (here leave the woods and walk through the prairies). Get out of the ravine by the paths that go up its top left side, and you will get to the Plleta del Perdigueret, a flat prairie where the water meanders at the bottom of the walls of the Perdiguero (2,380 m.). Then continue the ascent by the stony slope up to the Collada Ubaga (2,750 m.) where you will get after 3 and a half hours. Here the views widen, and you will see the large Ibón Blanco. The track now circumvents the Perdiguero by the rocky grounds of the slope (NE) and gets to the easy hill you will ascend to the top.

High Mountain Excursion. It is essential that you be accompanied by a High Mountain Guide (if you do not have previous experience), wear mountain boots and material such as crampons and ice axe. Not recommended for children.